# A Critical Assessment of the Brain Criticality Hypothesis

### 22 Apr 2026 v1

@@ III.1 Scope and limitations of the memory-based explanation
Of course, the best way to verify this behavior would be through experimentation. Specifically, since memory-induced long-range order relies upon two types of coupled variables, it would be ideal to study the degree of resource availability in the brain during normal spiking dynamics. Recent advances in genetically encoded sensors, [30] optical imaging techniques, [28, 8] and simultaneous metabolic-activity measurements [7] now make such experiments increasingly feasible. Furthermore, new research indicates that long-range order can even be “dragged” across layers in a multi-layered media. [48] Volumetric neural imaging techniques now exist [36, 12, 22] that could probe this behavior in the neocortex, which itself is organized into distinct layers. [18, 17, 25, 31]. We hope this opinion paper will spark interest in pursuing these experiments.
@@ 2. Possess long-range temporal correlations, and
Unlike Destexhe and Touboul’s counterexamples, memory-induced long-range order does satisfy each of these physically motivated requirements. First, the long-range order in these systems emerges precisely due to collective behavior in neural activity, which is induced by the presence of memory as provided by the resources. Additionally, since memory is necessarily present in this system, which inherently requires variables to correlate strongly over long durations, the existence of long-range correlations in time is also expected. This has been shown explicitly in the work of Sun et al. [43] by visualizing avalanche duration distributions, which also follow power-law decays. The first two points then follow directly from the presence of memory-induced long-range order.
The notion of “supporting information processing” is more nuanced. Several studies have argued that criticality optimizes properties like information transmission, [27, 4, 37, 41] storage, [27] computational power, [6] and more. However, these studies generally emphasize the coincidence of neuronal avalanches (and other hallmarks of long-range order) with heightened performance. Crucially, they do not require criticality, only power-law correlations. As such, these properties could well be optimized in systems with long-range order that are not necessarily critical.
Additionally, it is unclear whether the “optimal” operation of several brain processing functions is even biologically necessary. For instance, Wilting and Priesemann [46] have proposed that maximizing these processing functions simultaneously may not be required. Instead, perhaps performance must merely exceed some sufficient threshold for a given task. Beyond this point, it is feasible that the more relevant biological constraint would be minimizing energy consumption.
Lastly, recent computational work has indicated that long-range order alone may be sufficient for tasks requiring high computational performance. A new computational paradigm known as memcomputing [15, 16] has demonstrated that long-range order induced by memory, when applied to solving hard computational problems, can outperform state-of-the-art algorithms at various benchmarks, supporting the claim that memory-driven collective modes can be robust without fine tuning.
Altogether, we claim that much of the recent work intending to demonstrate that the brain operates at criticality does not always distinguish between criticality and long-range order or scale invariance. Without concrete evidence for the existence of a critical point and a better understanding of the phases directly above and below it, we instead suggest that the brain may naturally operate in an entire phase of (memory-induced) long-range order. The experimental fact that resources in the brain, which promote neural activity, are always necessarily present, naturally leads to such an explanation.
@@ V Glossary
Criticality Hypothesis - The hypothesis that the brain operates near a critical point of phase transition, and that at this point, several information processing functions operate optimally due to the collective interactions between neurons. [3]
Critical Point - An individual point separating two or more distinct phases or states. For example, in a ferromagnet, the Curie temperature marks the point where spontaneous magnetization disappears and the system becomes paramagnetic.
@@
Memory - A general property in physical systems in which the response of a system to some stimulus depends on the full history of its past dynamics. This is also known as time non-locality [^1] In this work, we will exclusively use the word “memory” in this context, rather than in the context of the storage or retrieval of information.
@@
Branching Process - A mathematical model which describes how information can be transferred in networks. These processes can be further categorized by their branching parameter, which counts how many neighboring nodes an individual node transfers, on average, information to. “Critical” branching processes have a branching parameter of one.

### 07 May 2026 v2

@@ III.3 How then can criticality be distinguished from a phase of long-range order?
Specifically, since MILRO relies upon two types of coupled variables, it would be ideal to study the degree of resource availability in the brain during normal spiking dynamics. Recent advances in genetically encoded sensors, [45] optical imaging techniques, [42, 12] and simultaneous metabolic-activity measurements [11] make such experiments increasingly feasible.
Past studies provide a clue as to what experimental techniques could facilitate this. In their 2005 paper, Haldeman and Beggs [32] demonstrated that avalanche statistics did change when their cortical samples were chemically altered to encourage excitation (more spiking) or inhibition (less spiking). As expected, this yielded avalanche size distributions that deviated somewhat from power-law decays, often called “super-critical” and “sub-critical,” respectively. This has been used as evidence that the central point is critical. [5]
However, a similar approach could be used in vitro to more thoroughly explore the phase diagram along the $\tau_{D}$ (memory timescale) direction. Traditional microelectrode arrays could be paired with optical devices to monitor vesicle recycling rates. [10, 58] Since vesicles play a key role in enabling neurotransmitters to propagate, thus gating the flow of electrical activity in neural systems, and are recycled over a much longer timescale than the spiking time of individual neurons, their recycling rate could act as an effective resource frequency. Once a baseline activity and resource timescale are established, 2-Deoxyglucose (2-DG), which inhibits glycolysis and thereby reduces glycolytic ATP production, could be then introduced. By slowing down ATP production, processes that rely upon ATP availability (like vesicle recovery) would slow down as well. Effectively, titrating 2-DG into neuronal cultures in vitro could slowly alter the effective resource timescale, thereby exploring a substantial line in the phase diagram. It is also possible that, in the near future, even more sophisticated tools could be applied similarly to measure neural activities [74] and intracellular ATP [45] in higher resolution.
Furthermore, new research indicates that long-range order can even be “dragged” across layers in multi-layered media. [90] Volumetric neural imaging techniques now exist [59, 19, 30] that could probe this behavior in the neocortex, which itself is organized into distinct layers. [25, 24, 33, 47]
Hopefully, this opinion paper will spark interest in pursuing these experiments.
@@ 2. Possess long-range temporal correlations, and
Memory-induced long-range order does satisfy each of these physically motivated requirements. First, the long-range order in these systems emerges precisely due to collective behavior in neural activity, which is induced by the presence of memory as provided by the resources. Additionally, since memory is necessarily present in this system, which inherently requires variables to correlate strongly over long durations, the existence of long-range correlations in time is also expected. This has been shown explicitly in the work of Sun et al. [75] by visualizing avalanche duration distributions, which also follow power-law decays. The first two points then follow directly from the definition of memory-induced long-range order. Of course, SOC could give rise to this behavior as well, but it requires further assumptions about the underlying system’s phase diagram than MILRO.
In regard to Beggs’ final point, recent computational work has indicated that long-range order alone may be sufficient for tasks requiring high computational performance. A new computational paradigm known as memcomputing [22, 23] has demonstrated that long-range order induced by memory, when applied to solving hard computational problems, can outperform state-of-the-art algorithms at various benchmarks, supporting the claim that memory-driven collective modes can be robust without fine tuning. Furthermore, recent studies have argued that the brain may exist in different dynamical states based on the external state or relevant computational task, [28, 13, 14] making the nature of the alleged critical state more ambiguous in these situations.
Nevertheless, much of the recent work intending to demonstrate that the brain operates at criticality does not always distinguish between criticality and long-range order or scale invariance. Studies displaying long-range correlated avalanche distributions, enhanced computational performance, or even a wide dynamic range [38, 69] can all be explained by either behavior. Without concrete evidence for the existence of a critical state, the authors argue that the brain may naturally operate in an entire phase of (memory-induced) long-range order. Other studies have argued how scale-invariance could arise in cortical dynamics away from criticality (via neutral drifts, [46] heavy-tailed connectivity, [70] stochasticity, [15] etc.), but to the authors’ knowledge, an analytical connection to time non-locality has not yet been made. The experimental fact that resources in the brain, which promote neural activity, act as a source of memory, naturally leads to such an explanation without requiring further evidence indicating criticality.
@@ V GLOSSARY
Criticality Hypothesis - The hypothesis that the brain operates near criticality, and that at this point, several information processing functions operate optimally due to the collective interactions between neurons. [7, 29, 79]
Criticality - A boundary of a continuous (second-order) phase transition separating two or more distinct phases or states. For example, a ferromagnet at the Curie temperature, where spontaneous magnetization disappears and the system becomes paramagnetic, is at criticality.
@@
Memory - A general property in physical systems in which the response of a system to some stimulus depends on the full history of its past dynamics. This is also known as time non-locality. In this work, the word “memory” is used exclusively in this context, rather than in the context of the storage or retrieval of information.
Self-Organized Criticality (SOC) - A process whereby systems are attracted dynamically in phase diagram to a point of continuous phase transition. This phenomenon requires a separation of timescales, including both slow external driving and fast relaxation. [2, 85]
@@
Branching Process - A mathematical model which describes how information can be transferred in networks. These processes can be further categorized by their branching parameter, which counts how many neighboring nodes an individual node transfers information to, on average. “Critical” branching processes have a branching parameter of one, although they are not necessarily associated with criticality from a phase diagram perspective.
Renormalization Group - A coarse-graining procedure which integrates out short-distance/high-energy degrees of freedom as part of the “RG flow.” This technique provides a more rigorous understanding of the phase diagram structure of a particular model.